OpenStudy (anonymous):

Find cos q if q is an angle in standard position and the point with coordinates (–12, 5) lies on the terminal side of the angle.

jimthompson5910 (jim_thompson5910):

here is the point (-12,5) 

jimthompson5910 (jim_thompson5910):

draw a right triangle like so  the cosine of angle q will be equal to the cosine of the lower angle on the right side in the triangle so cos(q) = adjacent/hypotenuse cos(q) = -12/13

OpenStudy (anonymous):

How did you find the degree on the unit circle?

OpenStudy (anonymous):

The point* not the degree.

jimthompson5910 (jim_thompson5910):

it's not on the unit circle if it was, then the hypotenuse would have been 1

jimthompson5910 (jim_thompson5910):

I'm just using the trig ratios in this case, cosine = adjacent/hypotenuse

jimthompson5910 (jim_thompson5910):

I guess I should have said that the 13 is from the fact that 5^2 + (-12)^2 = 13^2 by the pythagorean theorem

OpenStudy (anonymous):

oh so pretty much use the formula a^2+b^2=c^2?

jimthompson5910 (jim_thompson5910):

yes to find the missing length of a right triangle

OpenStudy (anonymous):

Ohhh okai thank you c:

jimthompson5910 (jim_thompson5910):

you're welcome